OpenStudy (anonymous):

You conduct 32 Bernoulli trials for which the probability of success is 0.18. What is the probability that you will obtain at most one success in your 32 trials? Write your answer correct to three decimal places.

OpenStudy (zarkon):

P(at most one success)=P(0 success or 1 success) =P(0 success)+P(1 success) \[=(1-.18)^{32}+32*(.18)*(1-.18)^{31}=.014\]